Assume the nominal rate of return is 5.31% and the real rate is 3.36%. Find the inflation rate of return using the exact formula.
Answer #1

The Fisher equation, which shows the exact relationship between nominal interest rates, real interest rates, and inflation is:

(1 + R) = (1 + r)(1 + h)

h = [(1 + .0531) / (1 + .0336)] – 1 = 0.0189 or 1.89%